When a potential difference of 146 V is applied to the plates of an air-filled parallel-plate capacitor, the plates carry a surface charge density of 2.30 ✕ 10−10 C/cm2. What is the spacing between the plates?
Answer #1

here,

the potential difference , V = 146 V

the surface charge density , sigma = Q /A

sigma = ( C * V )/A

sigma = ( A * e0 * V /d) /A

2.3 * 10^-10 = 8.85 * 10^-12 * 146 /d

d = 5.62 m

the seperation between the plates is 5.62 m
Answer #2

The 24 consonant sounds comprise six stops (plosives): p, b, t, d, k, g; the fricatives f, v, θ (as in thin), ð [eth] (as in then), s, z, [esh] (as in ship), Ʒ (as in pleasure), and h; two affricatives: t∫ (as in church) and dƷ (as the j in jam); the nasals m, n, ŋ (the sound that occurs at the end of words such as young); the lateral l; the postalveolar or retroflex r; and the semivowels j (often spelled y) and w. These remain fairly stable, but Inland Northern American differs from RP in two respects: (1) r following vowels is preserved in words such as door, flower, and harmony, whereas it is lost in RP; (2) t between vowels is voiced, so that metal and matter sound very much like British medal and madder, although the pronunciation of this t is softer and less aspirated, or breathy, than the d of British English.

answered by: Gamesome Goshawk